please simplify the radical.

5 square roots of a half - 2 square roots of an eighth.


Found 2 solutions by Alan3354, shymeerab:
Answer by Alan3354(69443)
Answer by shymeerab(21)

5(1/2)^(1/2)-2(8)^(1/2)
=5/(2^(1/2))-2(8)^(1/2)
=(5-2(2)^(1/2)(8)^(1/2))/2^(1/2)
=(5-2(16)^(1/2))/2^(1/2)
=((5-2*4)/2^(1/2))
=(5-8)/2^(1/2)
=-3/2^(1/2)